Portfolio theory

Given the following information about firm A and the market:δ

E (Rm)= 12 per cent                δm= 10 per cent ρA,M =0.95              Rf = 4 per cent         δA = 12 per cent

(i) Calculate the beta and the required rate of return for firm A.

(ii) Assume that an analyst, using fundamental analysis, estimates the expected return for firm A to be 15%. Is stock A overvalued or undervalued? Why?

(iii) Assume that an investor puts 70 per cent in the market portfolio and 40 per cent in the risk-free asset. What would the expected return and standard deviation of the portfolio be?
Answer #1

1.
Beta=0.95*12%/10%=1.14
Required return=4%+1.14*(12%-4%)=13.12%

2.
Undervalued as it is offering more return than is required

3.
Expected return=70%*12%+30%*12%=12.00%

Standard deviation=70%*10%=7%